OpenStudy (anonymous):

Simplify by distributing. -2(3a + 4).

jhonyy9 (jhonyy9):

-6a-8

OpenStudy (anonymous):

Step: -2(3a + 4) = -2(3a) + 4 (-2) = -6a - 8